# Exploring relaxation dynamics in warm dense plasmas by tailoring non-thermal electron distributions with a free electron laser

### v3

@@ I Introduction
Electron-phonon scattering is commonly assumed to be negligible in warm dense matter systems [11], and the other three contributions will dominate the relaxation process. Nevertheless, electron-phonon interactions could still be important in certain laser-produced plasmas, especially during plasma creation [12], and models can be constructed to account for such interactions, e.g. via the Boltzmann transport equation assuming near-equilibrium [13]. The contribution of e-n collisions, while not negligible, is generally believed to be well understood, and accepted cross-section models have been widely adopted in the literature [5, 11].
@@
There have been several attempts to correct the CL to allow it to be used at high densities. Some of the most recent attempts that still operate under the semi-classical Coulomb potential scattering (or Yukawa potential scattering) include applying Born’s approximation to cylindrical screening [15], or applying partial wave analysis [8, 17]. Beyond the semi-classical framework, there have also been quantum treatments, both within the framework of linear response theory [18, 19, 20, 21, 22] and by making use of other advanced methods [23, 24, 25, 26].
@@
A rich collection of theoretical approaches to the modelling of the CL, and of collisions more generally, has been developed over the past few decades. However, the validation of these models is severely hampered by the lack of experimental measurements of electron interaction dynamics and relaxation time scales in high-density systems. Very few attempts have been made to measure the electron relaxation time experimentally. Sun et al. reported a measured value of $\tau_{\rm e}\approx 1.7~{}{\rm fs}$ in fused silica [27] by measuring the plasma opacity in a pump-probe experiment. This method was developed by Pan et al. [28] to investigate the opacity dependence on electron density, but was limited to densities below that of a solid. Generally, there is little data for systems at densities around or exceeding that of a typical solid. It is worth stressing that determining plasma opacities is in and of itself a challenging problem in the warm-dense matter regime [29, 30], and the use of such data to infer collisional rates is often heavily model-dependent [31, 32].
@@
In recent years, x-ray Free Electron Lasers (XFELs) have proven capable of creating matter under extreme conditions of temperature, pressure, and density [33, 34, 35, 36]. Specifically, plasmas with a temperature of 100s of eV and remaining at exactly solid density have been created at LCLS by isochorical x-ray heating: during a typical $\sim 30$ fs XFEL pulse, the sample has no time to expand, and thus remaining at its original solid density, facilitating the study of transient warm dense plasmas in well-defined thermodynamics conditions [37]. Interestingly, the x-ray-matter interaction process by which samples are isochronally heated by an XFEL is via x-ray absorption and the generation of copious amounts of non-thermal electrons, either directly via photoionization or indirectly through Auger decay [38]. While the electrons subsequently collide and their distribution thermalizes, they also recombine into core holes, thus emitting photons that encode their distribution. By observing the emission spectrum from such non-thermal electrons, information regarding the electron distribution function and the characteristic relaxation time can thus be revealed.
@@
Here, we propose an experimental method to make use of this physical process to explore electron relaxation dynamics. Our approach focuses on tracking the evolution of non-thermal electrons within a broader electron distribution in a dense plasma on ultrafast time scales. This is achievable through the use of an XFEL, which can photoionize core electrons from atoms in a plasma, resulting in sustained and, importantly, well-defined non-thermal electron distributions. We present simulations conducted using the CCFLY collisional-radiative atomic kinetics code [39, 40, 36, 38]. Our results indicate that adjusting the XFEL pulse duration from sub-femtosecond to a few femtoseconds enables a measurement of the relaxation time scale of non-thermal electrons in the plasma, allowing us to glean insights into electron interaction processes on ultra-fast time scales and to assess the validity of the Coulomb Logarithm framework to describe dense plasma systems more generally.
@@
Figure 1: When a sample is driven by an intense x-ray pulse, K-shell photoionization and Auger decay create copious populations of non-thermal electrons at well-defined energies. These electrons thermalize collisionally with ions and other electrons in the system on electron equilibration time scales.
@@ II Method
When a sample is exposed to an XFEL beam with photon energy exceeding the $1s$ ionization energy (i.e., above the K-edge), the main absorption mechanism is K-shell photoionization (PI). In a low-Z element, the dominant relaxation of a $1s$ core hole takes place via Auger decay, which further excites an electron from a bound state into the continuum. These processes produce free electrons with well-defined energies, leading to non-thermal peaks in the electron energy distribution, as illustrated in Fig.  1. The non-thermal electrons created by this method thermalize rapidly yet are continuously replenished within the pulse duration of the driving XFEL beam. Therefore, there is a non-negligible probability for the non-thermal electrons to radiatively recombine into a core hole before the electron distribution fully equilibrates. Because the dominant radiative recombination channel will be into a $1s$ state, the emitted photons will have energies corresponding to the K-edge plus the energy of the free electron in the non-thermal distribution. As illustrated in Fig. 2, we can distinguish two main cases: recombination from PI electrons, with emitting photon energies similar to that of the driving x-ray pulse, and recombination of Auger-excited electrons, emitted at energies twice the Kα energy.
@@
Radiative recombination to the K shell can only take place in the presence of a K-shell core hole. These core holes are predominately produced through photon-ionization, with only a small fraction produced by collisional ionization due to the relatively low temperature compared with the K shell ionisation energy. Since these holes are majorly created by the XFEL beam, the X-ray pulse duration will determine the window within which emission can be observed. For a pulse length shorter than the relaxation time (typically below 1 fs), the majority of the emission will occur before significant thermalization can happen, leading to a strong non-thermal signature in the emission spectrum. In contrast, if the XFEL beam delivers the same number of photons but over a considerably longer pulse duration, a larger proportion of emission will take place from more thermalized electron distributions, leading to a weaker non-thermal signature in the emission spectrum. By varying the pulse duration to lie between these two limits, we can thus experimentally explore the electron equilibration process in a dense plasma. The ability of XFELs to deliver quasi-monochromatic x-ray pulses with durations ranging from sub-fs to several 10’s of fs thus makes them uniquely suited for probing electron relaxation times and, in particular, experimentally validating electron relaxation and transport models in dense plasmas.
II.1 Simulations
@@
We model the x-ray-matter interaction process using CCFLY [38], a non-LTE (local thermodynamic equilibrium) atomic-kinetics collisional-radiative code based on the popular SCFLY [40, 41] and FLYCHK [39] simulation suites previously used to model isochoric heating experiments at XFEL facilities [36]. Of importance to this work, CCFLY includes a Fokker-Planck solver for the free electron distribution function, which is coupled to the full non-LTE atomic kinetics framework. This allows us to model the time evolution of the electrons directly during the ultrafast x-ray isochoric heating process, and to compute a time-dependent emission spectrum selfconsistently. The CCFLY code tracks the atomic level populations as a function of time during the x-ray irradiation process, taking into account the full range of transitions between levels, including electron capture, collisional ionization and recombination, collisional transition, stimulated transitions, photo-ionization, inverse bremsstrahlung, and radiative recombination [39]. For processes involving free electrons, the transition rates make use of the non-thermal electron distribution function $f[E,t]$. The evolution of the electron distribution is determined by three components: the elastic collision term $\hat{\mathcal{O}}_{\rm e}[f]$, the inelastic collision term $\hat{\mathcal{O}}_{\rm i}[f]$, and a source term $S$ that accounts for ionization and recombination processes which do not conserve particle number. The model is zero-dimensional, so it does not account for any spatial extent of the system modelled. The charge state populations and the free electrons are thus modelled purely in terms of densities. In this model, the Boltzman-Vlasov equation for the evolution of the electron distribution is reduced to:
@@
The detailed forms of these terms can be found in refs. [42, 38], and references therein. Here, we only highlight that the thermalization is driven by both electron-electron (e-e) collisions and electron-ion (e-i) collisions. We will want to explore the importance of each of these terms to the overall equilibration process.
@@
While the x-ray isochoric heating process occurs on fs timescales, the equilibration timescale between electrons and ions is much longer, on the order of several ps. For the purposes of this work, we can, therefore, assume the ions are stationary and thus neglect the contribution of elastic e-i collisions. We adopt the cross-section model by Lotz [46] for collisional ionization and recombination, and that of Van Regenmorter [47] for collisional excitation and de-excitation. Radiative recombination and dielectronic recombination, though induced by e-i collisions, do not thermalize the electron distribution but only point-wisely subtract electrons at certain energies. They are thus captured by the source term $S$ in our simulation. Interestingly, e-e and e-i collisions thermalize the non-thermal electrons in two different ways. Typically, e-e collisions only change the energy of the electrons by a small amount per collision, but e-i collisions can remove a large amount of energy at a time. Intuitively, e-e collisions will thus tend to cause non-thermal peaks in the electron distribution to broaden while largely conserving the total number of electrons within it. In contrast, e-i collisions act to depopulate the non-thermal electron peaks but cause little broadening.
Figure 3: Evolution of the non-thermal electron distribution. (a) Non-thermal electrons are created via photoionization and Auger decay in the first fs of the pulse. (b) the electron distribution function thermalized to equilibrium after the x-ray pulse peaked at 1 fs.
The kinetic simulation computes the evolution of the atomic-level populations and of the electron distribution function $f[E,t]$. From these, the emissivity and opacity, and hence the time-dependent emission spectrum of the system, can be calculated. Integrating the emission over the duration of the X-ray pulse provides the observable that can, in principle, be measured experimentally. We note that Eq. (4) establishes a connection between the spectral emission from a system irradiated with x-rays of some given pulse duration and the collisional model setting the electron relaxation time scale. For example, in the Spitzer and Härm model [14], the relaxation time for e-e collisions is:
@@
where $k_{\rm B}$ is the Boltzmann constant, $T_{\rm e}$ is the effective electron temperature, and $n_{\rm e}$ is the electron density. The simulated emission spectrum is thus determined by the speed of relaxation given by Eq. (2), dependent on the Coulomb logarithm given by Eq. (4), which can be linked to a relaxation time via Eq. (5). In this work, we use Huba’s expression for the e-e Coulomb logarithm [48]. However, our Fokker-Planck treatment is compatible with other approaches and can be used to assess different Coulomb logarithm models.
@@ III Results
We first simulate the interaction of a 10 $\mu$m thick Al foil with an x-ray pulse with photon energy $\hbar\omega_{0}=2000~{}{\rm eV}$, pulse duration of 0.5 fs, and a peak intensity of $10^{19}$ Wcm-2 on target. The Al K-edge is located at $E_{\rm 1s}=1560~{}{\rm eV}$, so the photons carry sufficient energy to ionize the $1\rm s$ electrons. Our chosen focus spot sizes of 1 $\mu$m2 allow for the heating of solid density plasmas to electron temperatures of $\sim$ 70 eV.
[…]
Figure 5: Effect of XFEL pulse duration on the emission spectrum at constant pulse energy. a) The electron distribution when the XFEL intensity reaches its peak shows that for longer pulse durations, the electron distribution has been significantly thermalized. (b). Corresponding bound-free emission spectra show that the non-thermal feature vanishes when the pulse duration is long. Please note that we have used different normalisations for different pulse durations on the $y$-axis to provide a clearer comparison. The actual intensity of the emission spectra is similar across various pulse durations.
@@
III.1 Changing pulse duration

### v4

@@ I Introduction
Electron-phonon scattering is commonly assumed to be negligible in warm dense matter systems [11], and the other three contributions will dominate the relaxation process. Nevertheless, electron-phonon interactions could still be important in certain laser-produced plasmas, especially during plasma creation [12], and models can be constructed to account for such interactions, e.g. via the Boltzmann transport equation assuming near-thermal equilibrium [13]. The contribution of e-n collisions, while not negligible, is generally believed to be well understood, and accepted cross-section models have been widely adopted in the literature [5, 11].
@@
There have been several attempts to correct the CL to allow it to be used at high densities. Some of the most recent attempts that still operate under the semi-classical Coulomb potential scattering (or Yukawa potential scattering) include applying Born’s approximation to cylindrical screening [15], or applying partial wave analysis [8, 17]. There have also been several attempts that go beyond these CL approach, including the introduction of a quantum Landau-Fokker-Planck (qLFP) operator to the Boltzmann equation [18, 19, 20, 21], or using the semi-classical Boltzmann-Uehling-Ulenberg (BUU) equation [22, 23]. Further beyond the semi-classical framework, there have also been quantum treatments, both within the framework of linear response theory [24, 25, 26, 27, 28] and by making use of other advanced methods [29, 30, 31, 32].
@@
In contrast to the rich collection of theoretical development in the modelling of collision dynamics in the past few decades, the validation of these models is severely hampered by the lack of experimental measurements of electron interaction dynamics and relaxation timescales in high-density systems. Very few attempts have been made to measure the electron relaxation time experimentally. Sun et al. reported a measured value of $\tau_{\rm e}\approx 1.7~{}{\rm fs}$ in fused silica [33] by measuring the plasma opacity in a pump-probe experiment. This method was developed by Pan et al. [34] to investigate the opacity dependence on electron density, but was limited to densities below that of a solid. Generally, there is little data for systems at densities around or exceeding that of a typical solid. It is worth stressing that determining plasma opacities is in and of itself a challenging problem in the warm-dense matter regime [35, 36], and the use of such data to infer collisional rates is often heavily model-dependent [37, 38].
@@
In recent years, x-ray Free Electron Lasers (XFELs) have proven capable of creating matter under extreme conditions of temperature, pressure, and density [39, 40, 41, 42]. Specifically, plasmas with a temperature of 100s of eV and remaining at exactly solid density have been created at LCLS by isochorical x-ray heating: during a typical $\sim 30$ fs XFEL pulse, the sample has no time to expand, and thus remaining at its original solid density, facilitating the study of transient warm dense plasmas in well-defined thermodynamics conditions [43]. Interestingly, the x-ray-matter interaction process by which samples are isochronally heated by an XFEL is via x-ray absorption and the generation of copious amounts of non-thermal electrons, either directly via photoionization or indirectly through Auger decay [44]. While the electrons subsequently collide and their distribution thermalizes, they also recombine into core holes, thus emitting photons that encode their distribution. Similar techniques of utilizing free electron radiative recombination feature havs been well developed in electron-beam ion traps for measuring various important quantities, including crosssections and cooling rates [45, 46, 47], but only at densities far below that of a solid. They have also not been widely used in exploring the electron relaxation dynamics on an XFEL facility. In this work, we show that by observing the emission spectrum from such non-thermal electrons, information regarding the electron distribution function and the characteristic relaxation time can thus be revealed.
@@
Here, we propose an experimental method to make use of this physical process to explore electron relaxation dynamics. Our approach focuses on tracking the evolution of non-thermal electrons within a broader electron distribution in a dense plasma on ultrafast timescales. This is achievable through the use of an XFEL, which can photoionize core electrons from atoms in a plasma, resulting in sustained and, importantly, well-defined non-thermal electron distributions. We present simulations conducted using the CCFLY collisional-radiative atomic kinetics code integrated with a Fokker-Planck Collisional operator [48, 49, 42, 44]. Our results indicate that adjusting the XFEL pulse duration from sub-femtosecond to a few femtoseconds enables a measurement of the relaxation timescale of non-thermal electrons in the plasma. Notably, we have chosen to build our simulation based on the CL model to demonstrate the effectiveness of our method. This is not only because of the simplicity of CL but also due to it being well-known to a broad range of plasma physicists working on all density regimes. Nevertheless, all various types of collisional operators, including the abovementioned qLFP and BUU, are equally compatible and can be evaluated using the same experimental method proposed in this work. This allows us to glean insights into electron collision processes on ultra-fast timescales and to assess the validity of various assumptions used by the different collisional models in the warm dense regime.
@@
Figure 1: When a sample is driven by an intense x-ray pulse, K-shell photoionization and Auger decay create copious populations of non-thermal electrons at well-defined energies. These electrons thermalize collisionally with ions and other electrons in the system on electron thermalization timescales.
@@ II Method
When a sample is exposed to an XFEL beam with photon energy exceeding the $1s$ ionization energy (i.e., above the K-edge), the main absorption mechanism is K-shell photoionization (PI). In a low-Z element, the dominant relaxation of a $1s$ core hole takes place via Auger decay, which further excites an electron from a bound state into the continuum. These processes produce free electrons with well-defined energies, leading to non-thermal peaks in the electron energy distribution, as shown in Fig.  1. The non-thermal electrons created by this method thermalize rapidly yet are continuously replenished within the pulse duration of the driving XFEL beam. Therefore, there is a non-negligible probability for the non-thermal electrons to radiatively recombine into a core hole before the electron distribution fully thermalises. Because the dominant radiative recombination channel will be into a $1s$ state, the emitted photons will have energies corresponding to the K-edge plus the energy of the free electron in the non-thermal distribution. As illustrated in Fig. 2, we can distinguish two main cases: recombination from PI electrons, with emitting photon energies similar to that of the driving x-ray pulse, and recombination of Auger-excited electrons, emitted at energies twice the Kα energy.
@@
Radiative recombination to the K shell can only take place in the presence of a K-shell core hole. These core holes are predominately produced through photon-ionization, with only a small fraction produced by collisional ionization due to the relatively low temperature compared with the K shell ionisation energy. Since these holes are majorly created by the XFEL beam, the X-ray pulse duration will determine the window within which emission can be observed. For a pulse length shorter than the relaxation time (typically below 1 fs), the majority of the emission will occur before significant thermalization can happen, leading to a strong non-thermal signature in the emission spectrum. In contrast, if the XFEL beam delivers the same number of photons but over a considerably longer pulse duration, a larger proportion of emission will take place from more thermalized electron distributions, leading to a weaker non-thermal signature in the emission spectrum. By varying the pulse duration to lie between these two limits, we can thus experimentally explore the electron thermalization process in a dense plasma. The ability of XFELs to deliver quasi-monochromatic x-ray pulses with durations ranging from sub-fs to several 10’s of fs thus makes them uniquely suited for probing electron relaxation times and, in particular, experimentally validating electron relaxation and transport models in dense plasmas.
II.1 The collisional radiative model
@@
We model the x-ray-matter interaction process using CCFLY [44], a non-LTE (local thermodynamic equilibrium) atomic-kinetics collisional-radiative code based on the popular SCFLY [49, 50] and FLYCHK [48] simulation suites previously used to model isochoric heating experiments at XFEL facilities [42]. Of importance to this work, CCFLY includes a Fokker-Planck solver for the free electron distribution function, which is coupled to the full non-LTE atomic kinetics framework. This allows us to model the time evolution of the electrons directly during the ultrafast x-ray isochoric heating process, and to compute a time-dependent emission spectrum selfconsistently. The CCFLY code tracks the atomic level populations as a function of time during the x-ray irradiation process. This is achieved by self-consistently solving the rate equations:
$\frac{\mathrm{d}n_{i}}{\mathrm{~{}d}t}=\sum_{j}M_{i,j}n_{j},$ (1)
where $n_{i}$ is the population of the $i^{\rm th}$ atomic state, and the matrix elements are determined by the transition rates $R_{i\rightarrow j}$ through:
$M_{i,j}=\left\{\begin{array}[]{cc}R_{j\rightarrow i}&(j\neq i)\\ -\sum_{k}R_{i\rightarrow k}&(j=i)\end{array}\right..$ (2)
For simplicity purposes, we chose to use the super configuration atomic states with maximum principle quantum number $n^{*}=4$, so all atomic states are expressed in the form of ${\rm K}^{a}{\rm L}^{b}{\rm M}^{c}{\rm N}^{d}$, where $a,b,c,d$ denotes the number of electrons on each shell. The transition rates $R_{j\rightarrow i}$ take into account the full range of transitions between these two levels, including electron capture, collisional ionization and recombination, collisional transition, stimulated transitions, photo-ionization, inverse bremsstrahlung, and radiative recombination [48]. We highlight that processes which involve free electrons, for example, collisional ionization (CI), depend on the free electron distribution:
$R_{j\rightarrow i,{\rm CI}}=\int_{\vec{v}}v\varrho[\vec{v}]\sigma_{j\rightarrow i,{\rm CI}}[\vec{v}]\mathrm{~{}d}v_{x}\mathrm{d}v_{y}\mathrm{d}v_{Z}.$ (3)
II.2 Electron relaxation
For processes involving free electrons, the transition rates make use of the non-thermal electron distribution function $f[E,t]$. The evolution of the electron distribution is determined by three components: the elastic collision term $\hat{\mathcal{O}}_{\rm e}[f]$, the inelastic collision term $\hat{\mathcal{O}}_{\rm i}[f]$, and a source term $S$ that accounts for ionization and recombination processes which do not conserve particle number. The model is zero-dimensional, so it does not account for any spatial extent of the system modelled. The charge state populations and the free electrons are thus modelled purely in terms of densities. In this model, the Boltzman-Vlasov equation for the evolution of the electron distribution is reduced to:
@@ II.2 Electron relaxation
The detailed forms of these terms can be found in refs. [51, 44], and references therein. Here, we only highlight that the thermalization is driven by both electron-electron (e-e) collisions and electron-ion (e-i) collisions. We will want to explore the importance of each of these terms to the overall thermalization process.
@@
While the x-ray isochoric heating process occurs on fs timescales, the thermalization timescale between electrons and ions is much longer, on the order of several ps. For the purposes of this work, we can, therefore, assume the ions are stationary and thus neglect the contribution of elastic e-i collisions. We adopt the cross-section model by Lotz [55] for collisional ionization and recombination, and that of Van Regenmorter [56] for collisional excitation and de-excitation. Radiative recombination and dielectronic recombination, though induced by e-i collisions, do not thermalize the electron distribution but only point-wisely subtract electrons at certain energies. They are thus captured by the source term $S$ in our simulation. Interestingly, e-e and e-i collisions thermalize the non-thermal electrons in two different ways. Typically, e-e collisions only change the energy of the electrons by a small amount per collision, but e-i collisions can remove a large amount of energy at a time. Intuitively, e-e collisions will thus tend to cause non-thermal peaks in the electron distribution to broaden while largely conserving the total number of electrons within it. In contrast, e-i collisions act to depopulate the non-thermal electron peaks but cause little broadening.
Figure 3: Evolution of the non-thermal electron distribution. (a) Non-thermal electrons are created via photoionization and Auger decay in the first fs of the pulse. (b) the electron distribution function thermalized to equilibrium after the x-ray pulse peaked at 1 fs. Note that the electron density is not a constant as electrons are removed by recombination and collisional processes. We discuss the effect of these processes in Sec III/B
We note that Eq. (7) establishes a connection between the spectral emission from a system irradiated with x-rays of some given pulse duration and the collisional model setting the electron relaxation timescale. For example, in the Spitzer and Härm model [14], the relaxation time for e-e collisions is:
@@
where $k_{\rm B}$ is the Boltzmann constant, $T_{\rm e}$ is the effective electron temperature, and $n_{\rm e}$ is the electron density. The simulated emission spectrum is thus determined by the speed of relaxation given by Eq. (5), dependent on the CL given by Eq. (7), which can be linked to a relaxation time via Eq. (8). In this work, we use Huba’s expression for the e-e Coulomb logarithm [57]. However, our Fokker-Planck treatment is compatible with other approaches and can be used to assess different collision models.
We note that this relaxation time can only be used to quantify the collision frequency in the thermal bulk, and the timescale for non-thermal electrons reaching a thermalized distribution can be different. In general, electrons will higher energy will thermalize slower. However, the functional nature of $W[\varrho,v]$ on the right-hand side of Eq. 5 indicates no simple way to define such a quantity consistently as one would usually do in relaxation time approximation ($\frac{\partial f[E]}{\partial t}\approx\frac{f[E]-f_{0}[E]}{\tau}$, where $f_{0}$ is the thermal distribution), even if an explicit dependence on electron energy is included. This is because the rate evaluated at a certain energy $E$ is not only affected by the pointwise value $f[E]$ but rather the whole distribution. Nevertheless, under the assumption that the non-thermal fraction remains small, the relaxation timescale given by Eq. 5 still indicates the overall collision rate, which is the key quantity of interest that captures the dependence on collision models, and can be evaluated by the method we proposed in this worl
@@ III Results
We first simulate the interaction of a 10 $\mu$m thick Al foil with an x-ray pulse with photon energy $\hbar\omega_{0}=2000~{}{\rm eV}$, pulse duration of 0.5 fs, and a peak intensity of $10^{19}$ Wcm-2 on target. In this simulation, we assumed a Gaussian-shaped beam profile both in the time and frequency domain. We have chosen a bandwidth of 0.8% $\times$ incident photon energy, in this case, 16 eV, which is achievable at XFEL facilities under the XLEAP operation mode [58] for sub-fs pulse duration. We note that the intrinsic Fourier limit of the X-ray pulse becomes larger for shorter pulse operations. However, this only becomes comparable to our 10 eV setup when the pulse duration is reduced to less than $4\ln{2}\frac{\hbar}{10~{}{\rm eV}}\approx 0.1~{}{\rm fs}$, validating our assumption of XFEL bandwidth. The Al K-edge is located at $E_{\rm 1s}=1560~{}{\rm eV}$, so the photons carry sufficient energy to ionize the $1\rm s$ electrons. Our chosen focus spot sizes of 1 $\mu$m2 allow for the heating of solid density plasmas to electron temperatures of $\sim$ 70 eV. This equates to having 50 uJ energy, or $\sim 10^{11}$ photons in the beam. Both the pulse energy and the focusing are readily achievable at various endstations in conventional XFEl facilities like MEC at LCLS and HED at EuXFEL. Better focusing may be achieved by the nanofocusing technique [59], whilst we find defocusing more interesting as this would help to emphasize low-temperature relaxation dynamics.
@@
[…]
Figure 5: The effect of the XFEL pulse duration on the emission spectrum at constant pulse energy. a) The electron distribution when the XFEL intensity reaches its peak is drawn to show that for longer pulse durations, the electron distribution has been significantly thermalized. (b). Corresponding bound-free emission spectra show that the non-thermal feature vanishes when the pulse duration is long. Emission spectra from pulses with different durations have been offset on the $y$-axis to provide a clearer comparison. The actual overall intensities of all these emission spectra are similar across various pulse durations.
@@
III.1 Changing the pulse duration